# Human-Robot Interaction Conversational User Enjoyment Scale (HRI CUES)

### v2

@@ Abstract
Understanding user enjoyment is crucial in human-robot interaction (HRI), as it can impact interaction quality and influence user acceptance and long-term engagement with robots, particularly in the context of conversations with social robots. However, current assessment methods rely solely on self-reported questionnaires, failing to capture interaction dynamics. This work introduces the Human-Robot Interaction Conversational User Enjoyment Scale (HRI CUES), a novel scale for assessing user enjoyment from an external perspective during conversations with a robot. Developed through rigorous evaluations and discussions of three annotators with relevant expertise, the scale provides a structured framework for assessing enjoyment in each conversation exchange (turn) alongside overall interaction levels. It aims to complement self-reported enjoyment from users and holds the potential for autonomously identifying user enjoyment in real-time HRI. The scale was validated on 25 older adults’ open-domain dialogue with a companion robot that was powered by a large language model for conversations, corresponding to 174 minutes of data, showing moderate to good alignment. The dataset is available online111HRI CUES Dataset (anonymized transcripts, annotation scores, and self-reported user perceptions): https://paperswithcode.com/dataset/hri-cues-dataset. Additionally, the study offers insights into understanding the nuances and challenges of assessing user enjoyment in robot interactions, and provides guidelines on applying the scale to other domains.
@@
User enjoyment, referring to the user’s subjective perception and experience of the enjoyment of interaction, is an important indicator of acceptance of robots and willingness to engage with them over time [1]. Particularly in the context of conversational agents or companion robots, where the primary goal often revolves around providing emotional support or companionship, enjoyment serves as a vital metric for evaluating the effectiveness of such systems. Therefore, developing reliable and efficient methods for measuring user enjoyment in Human-Robot Interaction (HRI) scenarios is essential for designing and improving future generations of robots.
@@
User enjoyment is closely linked to the intention to use robots, particularly among older adults [1]. Conversational companion robots are often developed to provide social or emotional support to older adults in a home or care home environment [2]. Prior studies in HRI explored older adults’ acceptance, use, and interaction with robots, showing that older adults often display challenges in interacting with a conversational agent, such as the lack of conversational responses and difficulties in hearing and understanding the voice interaction [3], consequently attributing a low level of social acceptance to the robot [4]. Designing robots with their individual needs and preferences in mind by involving them in the design process with participatory design research techniques, such as focus groups, interviews, and iterative developments based on their feedback, could potentially help alleviate these interaction challenges [5, 3, 6]. The recent introduction of Large Language Models (LLMs) has enabled the development of companion robots equipped with social capabilities, eliminating the need for Wizard of Oz, which was the common approach in conversational HRI studies (e.g., [3, 4]), and the inherent human influence that hinders the construction of robots capable of autonomously mitigating errors [7, 8]. Recent studies applied LLMs to conversational robots in various domains, including therapy [9], service [10], and care for older adults [11, 12], which demonstrate their potential and limitations in diverse contexts that lead to enjoyable or unpleasant experiences, further showing the importance of detecting user enjoyment during conversations with robots.
@@
Sustaining enjoyment, especially in daily encounters such as for companion robots, is a challenging task yet to be solved. User interest and engagement may fluctuate within day-to-day interactions, but also within the interaction itself, based on the robot’s performance in conversation flow, content, and contextual memory, which may affect user enjoyment. If user enjoyment can be detected autonomously in the conversation, preventive measures can be taken to improve the interaction where necessary, such as changing the conversation topic. However, despite the numerous studies investigating user engagement in HRI [13], measure of user enjoyment is limited to self-reports from users. Not only can self-reports be unreliable due to demand characteristics, self-presentation, or Hawthorne effect due to conformity to perceived norms or researcher expectations [14], but they represent an overall feedback of the interaction rather than an instantaneous measure throughout. While affect recognition systems can detect laughter and smiles [15], enjoyment is a complex feeling that can be conveyed through other multimodal cues (see Section IV-C). Even in interpersonal communication, enjoyment has been analyzed from an external perspective only within the context of marriage [16]. Thus, there is no scale or an automatic system for assessing user enjoyment in conversations with a robot, the former being required to develop the latter.
@@
This work contributes with the Human-Robot Interaction Conversational User Enjoyment Scale (HRI CUES), illustrated in Fig. 1, which is a novel scale for assessing enjoyment in conversations with robots from an external (third-party) perspective, deriving from videos of 28 older adults’ open-domain dialogue with a companion robot using an LLM. The scale is developed through rigorous annotator evaluations and discussions, and provides a structured framework for evaluating enjoyment in conversations with robots. The scale seeks to offer an additional means of assessing user enjoyment in HRI by considering fine-grained conversation exchange levels (i.e., turn-by-turn) and the overall interaction level, which complements self-reported enjoyment from users, with future potential use for autonomously identifying enjoyment in real-time HRI. In addition, by providing a detailed exploration of the instances of annotators’ concordance and divergence, based on turn-by-turn analysis of enjoyment, in addition to the underlying reasons for discrepancies between users’ self-reported enjoyment ratings based on metrics typically used in HRI studies, the study offers invaluable insights for understanding the nuances and challenges of assessing user enjoyment in interactions with robots. Deriving from these challenges, a step-by-step guideline is offered for future HRI researchers to adapt the user enjoyment scale to other application domains.
@@ II-A Defining enjoyment
A popular definition of enjoyment is being in the state of flow [17]. Flow is defined as the optimal experience, which provides a deep sense of enjoyment. It happens when an individual is fully engaged in a task that provides an optimal amount of challenge and engagement. Flow is characterized by a set of factors, such as a fading sense of ‘self’, a sense that duration is altered, and deep and effortless involvement in the task. The theory states that enjoyment is not obtained in a relaxed state, that it is necessary to be challenged, and links repeated experiences of flow to mastery of a skill.
@@
While both flow and true fun emphasize high engagement through optimal challenge, other theories recognize enjoyment in less intense states. For instance, flow-like states can be differentiated from the overall positive valence of an experience [19]. This relates to the circumplex model of emotion [20], which features arousal (low to high) on one axis and pleasure or valence (negative to positive) on the other axis [21]. Certain theories of enjoyment focus on high valence values that can contain lower-engagement positive emotions (e.g., content or calm), while others prioritize high-arousal states (e.g., excitement) [19]. The ‘happy’ emotion in the circumplex model reflects a balance of high arousal and positive valence. Our work aligns with the theory of true fun in seeking both higher levels of arousal and valence, while incorporating other elements for classifying lower levels of enjoyment. Casual, open-domain conversations between older adults and robots may involve aspects of both arousal and positive emotions. As such, both factors will be relevant to our holistic model of enjoyment.
@@ II-B Assessing enjoyment
Enjoyment is generally evaluated through self-reported questionnaires tailored to the specific application domain. For instance, the Quality of Life Enjoyment and Satisfaction Questionnaire (Q-LES-Q) [22] and Physical Activity Enjoyment Scale (PACES) [23] are used in healthcare applications. In Human-Computer Interaction (HCI) and HRI research, enjoyment is not often the primary focus for evaluating user perceptions, but is typically included as part of a more comprehensive model [24]. Enjoyment frequently appears as a self-reported measure, either as a construct within established models like the Unified Theory of Acceptance and Use of Technology (UTAUT) [25], or as single-item measures in custom questionnaires (e.g.,“Did you feel fun?” [26], “Was playing with the robot enjoyable/not enjoyable?” [27]). Enjoyment was found to be highly correlated with ‘satisfying’, ‘entertaining’, ‘exciting’,‘fun’, and ‘interesting’ in HRI [28]. Technology Acceptance Model (TAM) [29] was also adapted to HRI by incorporating measures of affect and cognition to improve its accuracy in explaining technology adoption, and this adaptation included questions about perceived enjoyment [30].
@@
While user enjoyment is commonly measured through self-reporting in HRI, it has several limitations, such as conforming to perceived norms or researcher expectations or the (in)ability to recall the events and report correctly from memory [33, 14]. In addition, it is often desirable to estimate what a user is feeling by assessing it from an external perspective when self-reporting is not possible or the goal is to automate behavior at the dialogue exchange level during interactions in real-time. External assessments and self-reporting are not mutually exclusive and can instead complement each other. The results of one method can even be used to validate the other. While prior research used smiles and laughter for automatic classification of user enjoyment (within the context of story-telling) [15], these signals can be contradictory, ambiguous and highly context dependent [34, 35], and enjoyment is often expressed through other multimodal cues.
@@ II-C Assessing enjoyment in conversations
User satisfaction, which correlates significantly with user enjoyment in conversations [36, 37], is typically evaluated in relation to a task [38, 39]. For instance, the Paradigm for Dialogue System Evaluation (PARADISE) [38] is a framework for evaluating user satisfaction in dialogue systems, based on self-reported satisfaction on a dialogue level and is influenced by other metrics, such as task success in travel booking and accessing emails. Similarly, interaction quality, which evaluates user satisfaction from an external perspective on the exchange level, was analyzed by three annotators in bus schedule inquiries with chatbots over phone calls from a data corpus of 200 dialogues and a lab study with 38 subjects [39]. An autonomous system was developed based on their ratings, which correlated highly with them, but not with users’ self-reported satisfaction scores, which was attributed to the subjectivity of the measure and variability in user perceptions.
User satisfaction has also been measured in the text domain with an annotation protocol similar to our study, based on a dataset of 1000 dialogues between 50 users and a chatbot on attentive listening and conversations about animals [40]. Two annotators were recruited and an annotator alignment session was conducted. The annotators were requested to go through the conversation exchanges once, without going back or looking at the history. The annotators used three metrics of user satisfaction: ‘smoothness of the conversation’, ‘closeness perceived by the user towards the system’, and ‘willingness to continue the conversation’, rated from 1 to 7. However, no agreement was found between the annotators, even after changing the granularity of the scale to two levels, low and high, showing the complexity of evaluating a subjective measure from an external perspective. The study focused on developing an algorithm to measure user satisfaction.
A similar study, also exploring user satisfaction, was conducted to develop a multimodal model based on annotated data [41]. The data corpus consisted of conversations of 60 participants with a virtual agent that was controlled through Wizard of Oz. The wizard annotated the user satisfaction from their perspective after each dialogue exchange, and the user and wizard both provided the satisfaction level of the user at the end of the conversation. In addition, five annotators were recruited to annotate the dialogues from an external perspective. All annotators gave each exchange in the conversation a score in the following metrics: topic continuance (1: ‘strongly change the topic’ to 7: ‘strongly continue the topic’), external sentiment (1: ‘the participants seemed bored with the dialogue’ to 7: ‘participants seemed to enjoy the dialogue’), and self-sentiment (1: ‘want to stop talking/confused about the system utterances’ to 7: ‘enjoy talking/satisfied with the talk’). Notably, the two latter were labeled enjoyment on the upper end of the scale. They received a good level of agreement between the annotators on the exchange level. To annotate the dialogue level, a questionnaire was used with 18 items that were designed to represent three labels: ‘coordinateness’, ‘awkwardness’, and ‘friendliness’. Based on these criteria, the developed multimodal model has outperformed the annotators’ in evaluating user satisfaction at the overall interaction level.
@@
Our user enjoyment scale was developed based on the work of Reimnitz and Rauer [16], which is the only scale that specifically evaluates user enjoyment in conversations from an external perspective. The study assessed the enjoyment in conversations between 64 married couples and compared that to each spouse’s marital happiness. For measuring enjoyment, they developed a scale for observational coding that took into consideration affective signs and the tone of the interaction. The scale ranged from 1 (very low enjoyment) to 7 (very high enjoyment), with 3 as a neutral anchor. Two annotators took into account both affective signs (e.g., mutuality of the interaction, tone of voice, consistent mutual gaze, facial expressions, physical touching, body language) and the tone of the interaction (i.e., neutral, enthusiastic, and delightful) when rating enjoyment. The annotators had good to moderate agreement, with intraclass correlation (ICC) on 20% of the interactions. The study found that couples who displayed high enjoyment in their conversations also reported having a happier marriage. This aligns with prior research in human relationships, which found that mutually enjoyable behavior leads to increased intimacy, trust, security, and satisfaction in long-term relationships [42], signaling that enjoyment could be highly influential in achieving long-term HRIs.
@@
Similar to this scale, our study analyzes dyadic conversations outside of task-oriented settings, but with a focus on evaluating enjoyment in conversations with autonomous robots. Although prior research described in this section evaluated autonomous conversational systems using user satisfaction metrics, which touch upon aspects of enjoyment, these metrics are often more focused on task-specific aspects. In contrast, our study emphasizes understanding enjoyment within daily (open-domain) conversations with robots, based on older adults’ interactions with a companion robot. To the best of our knowledge, no prior study measures enjoyment in open-domain conversations with robots from an external perspective that take into account multimodal aspects of HRI. Our work seeks to bridge this gap by proposing a scale that not only emphasizes the importance and complexity of user enjoyment, but also provides a methodological framework to assess it from an external perspective in other application domains of HRI. The scale aims to provide an additional tool to evaluate enjoyment complementary to self-reported measures from users within conversation exchange and overall interaction levels, with the potential to be used for autonomous systems to adapt the conversations on the fly.
III Data Collection
This work aims to offer an additional tool for assessing user enjoyment in conversational HRI. Our research seeks to evaluate perceived enjoyment within conversation exchanges and interaction levels from an external perspective, to complement self-reported questionnaires and provide a holistic view of the interaction as it progresses. To achieve this, we embarked on developing a user enjoyment scale and validating it within the context of an HRI scenario. As outlined in Section I, achieving and maintaining user enjoyment is important for engaging users and encouraging continued interactions with robots, especially in daily encounters. This becomes particularly prominent for companion robots for older adults that aim to provide social and emotional support to mitigate loneliness in their daily lives. Having daily conversations spanning a wide range of topics, i.e., open-domain dialogue, plays a pivotal role in achieving this, as it enables more engaging and fulfilling conversations that cater to the user’s various emotional and cognitive needs. While this has been challenging to achieve in the past with conversational robots [43], recent advancements in LLMs enable these capabilities, which make them suitable architectures for companion robots. In addition, these robots should be tailored to older adults’ unique needs and preferences to meet their expectations and provide them with enjoyable conversations. Thus, it is crucial to iteratively involve older adults in the design process to both learn their needs and effectively assess the developed systems.
Thus, this work builds upon the data from our prior work on the participatory design development of an autonomous companion robot that integrates an LLM for conversations with older adults, as described in [12], to build the user enjoyment scale for conversational HRI and evaluate enjoyment. Initially, preliminary interviews were conducted with 6 Swedish-speaking older adults who talked to the robot individually for 4-13 min. Based on the feedback from the initial study, the robot architecture was improved to eliminate initial challenges. Following that, four design workshops were conducted with 28 older adults who talked to the robot for 7 min with surveys and interviews that followed. This section summarizes the robot architecture used and the study details, described in detail in [12].

### v3

@@ Abstract
Understanding user enjoyment is crucial in human-robot interaction (HRI), as it can impact interaction quality and influence user acceptance and long-term engagement with robots, particularly in the context of conversations with social robots. However, current assessment methods rely solely on self-reported questionnaires, failing to capture interaction dynamics. This work introduces the Human-Robot Interaction Conversational User Enjoyment Scale (HRI CUES), a novel 5-point scale to assess user enjoyment from an external perspective (e.g., by an annotator) for conversations with a robot. The scale was developed through rigorous evaluations and discussions among three annotators with relevant expertise, using open-domain conversations with a companion robot that was powered by a large language model, and was applied to each conversation exchange (i.e., a robot–participant turn pair) alongside overall interaction. It was evaluated on 25 older adults’ interactions with the companion robot, corresponding to 174 minutes of data, showing moderate to good alignment between annotators. Although the scale was developed and tested in the context of older adult interactions with a robot, its basis in general and non-task-specific indicators of enjoyment supports its broader applicability. The study further offers insights into understanding the nuances and challenges of assessing user enjoyment in robot interactions, and provides guidelines on applying the scale to other domains and populations. The dataset is available online111HRI CUES Dataset (anonymized transcripts, annotation scores, and self-reported user perceptions): https://doi.org/10.5281/zenodo.12588810.
@@
User enjoyment, referring to the user’s subjective perception and experience of the enjoyment of interaction, is an important indicator of acceptance of robots and willingness to engage with them over time [1]. Particularly in the context of conversational agents or companion robots, where the primary goal often revolves around providing emotional support or companionship, enjoyment serves as a vital metric for evaluating the effectiveness of such systems. Therefore, developing reliable and efficient methods to measure user enjoyment in Human-Robot Interaction (HRI) scenarios is essential for designing and improving future generations of robots.
@@
User enjoyment is closely linked to the intention to use robots, particularly among older adults [1]. Conversational companion robots are often developed to provide social or emotional support to older adults in a home or care home environment [2]. Prior studies in HRI explored older adults’ acceptance, use, and interaction with robots, showing that older adults experience difficulties in interacting with a conversational agent, such as in hearing and understanding the system, and responding to it [3], consequently attributing a low level of social acceptance to the robot [4]. Some of these challenges emerged from the limitations of earlier systems and approaches, which relied heavily on scripted and non-adaptive interactions, underscoring the need for more autonomous and socially capable robots. The recent introduction of Large Language Models (LLMs) has enabled the development of companion robots equipped with social capabilities, eliminating the need for scripted interactions or Wizard of Oz, which was the common approach in conversational HRI studies (e.g., [3, 4]), and the inherent human influence that hinders the construction of robots capable of autonomously mitigating errors [5, 6]. Recent studies applied LLMs to conversational robots in various domains, including therapy [7], service [8], and care for older adults [9, 10], which demonstrate their potential and limitations in diverse contexts that lead to enjoyable or unpleasant experiences, further showing the importance of detecting user enjoyment during conversations with robots.
@@
Sustaining enjoyment, especially in daily encounters such as for companion robots, is a challenging task yet to be solved. User engagement and satisfaction may fluctuate within day-to-day interactions, but also within the interaction itself, based on the robot’s performance in conversation flow, content, and contextual memory, which may affect user enjoyment. However, relying solely on user engagement or task satisfaction as an indicator of enjoyment can be misleading, as these are distinct concepts that may occur simultaneously but can also be mutually exclusive, that is, one may exist without the other depending on the context of the interaction. For instance, a user can be engaged in an argument or be satisfied with the agent completing a task, but not necessarily enjoy either interaction. Despite several studies detecting user engagement [11] or satisfaction [12, 13] with automatic measures, measure of user enjoyment is limited to self-reports from users [14, 15]. Not only can self-reports be unreliable due to demand characteristics, self-presentation, or the Hawthorne effect (stemming from conformity to perceived norms or researcher expectations [16]), but they represent overall feedback of the interaction rather than a continuous, moment-by-moment measure. While affect recognition systems can detect laughter and smiles [17], enjoyment is a complex feeling that can be conveyed through other multimodal cues (see Section IV-D). Even in interpersonal communication, enjoyment has been analyzed from an external perspective mainly within the context of marriage [18] or computer entertainment [19]. Thus, there is no scale or an automatic system for assessing user enjoyment in conversations with a robot, the former being required to develop the latter.
@@
This work contributes with the Human-Robot Interaction Conversational User Enjoyment Scale (HRI CUES), illustrated in Fig. 1, which is a novel 5-point scale for assessing enjoyment in conversations with robots from an external (third-party) perspective. The scale was developed through rigorous annotator discussions based on individual open-domain conversations of three older adults with a companion robot using an LLM, and was tested on 25 older adults’ conversations with the same robot, achieving moderate to good annotator alignment. This 1-item scale is not intended to be psychometric; rather, it serves as a methodological tool for annotators to assess user enjoyment construct in HRI, by considering fine-grained conversation exchange levels (i.e., turn-by-turn) and the overall interaction level. The scale also holds potential for autonomously identifying enjoyment in real-time HRI [20, 21]. HRI CUES is presented in a generic form, which has to be adapted to each specific use case (as described in Section VI-C). In the particular case described in this paper, open-ended conversations with older adults in Sweden were used (described in Section III) both for the empirical backbone in the development of HRI CUES and for proof of concept. In addition, by providing a detailed exploration of the instances of annotators’ concordance and divergence, based on turn-by-turn analysis of enjoyment, in addition to the underlying reasons for discrepancies between users’ self-reported enjoyment ratings based on metrics typically used in HRI studies, the study offers invaluable insights for understanding the nuances and challenges of assessing user enjoyment in interactions with robots.
@@ II-A Defining enjoyment
Evaluating the user’s subjective experience is extremely important in Human-Computer Interaction (HCI) and HRI, with ‘enjoyment’ being a frequently assessed metric, e.g., [1, 22, 23, 15]. In this paper, we conceptualize enjoyment as a felt experience whose nature is primarily determined by two key characteristics: first, its hedonic quality, indicating the extent to which it is perceived as pleasant or unpleasant; and second, its intensity level, describing how activated or subdued the experience is. This approach allows us to consider not only intense positive states like ‘flow’ and relaxed positive feelings, but also to acknowledge and measure experiences involving negative hedonic qualities. Such a comprehensive view of enjoyment is vital for a holistic assessment of user affect in nuanced interactions.
A previous common definition of enjoyment is being in the state of flow [24]. Flow is defined as the optimal experience, which provides a deep sense of enjoyment. It happens when an individual is fully engaged in a task that provides an optimal amount of challenge. Flow is characterized by a set of factors, such as a fading sense of ‘self’, a sense that duration is altered, and deep and effortless involvement in the task. The theory states that enjoyment is not obtained in a relaxed state, that it is necessary to be challenged, and links repeated experiences of flow to mastery of a skill.
@@
User enjoyment can align with engagement as evidenced in the theory of flow where a sense of deep focus is highlighted. However, a person might be deeply engaged in an argument, being focused, responsive, and mentally invested, yet not enjoy the interaction at all [26]. Conversely, someone might enjoy the atmosphere of a group conversation without being particularly engaged with the robot or any individual speaker. Similarly, user satisfaction is related to enjoyment and is defined as “freedom from discomfort and positive attitudes towards the use of the product” [27]. However, user satisfaction is an evaluation of the user’s experience of a system performing a task, focusing on the task itself. For instance, a user might report satisfaction with a robot’s performance because it met their expectations or completed a task successfully, even if the interaction itself was not enjoyable.
Other theories recognize enjoyment in less intense states without focusing on the difficulty of the task. For instance, flow-like states can be differentiated from the overall positive valence of an experience [28]. This relates to the circumplex model of emotion [29], which features arousal (low to high) on one axis and pleasure or valence (negative to positive) on the other axis [30]. Certain theories of enjoyment focus on high valence values that can contain lower-engagement positive emotions (e.g., content or calm), while others prioritize high-arousal states (e.g., excitement) [28]. The ‘happy’ emotion in the circumplex model reflects a balance of high arousal and positive valence.
Our work aims to capture both higher levels of arousal and, especially, valence, while incorporating other elements for classifying lower levels of enjoyment. Casual, open-domain conversations between older adults and robots may involve aspects of both arousal and a range of emotions. As such, both dimensions are relevant to our holistic model of enjoyment.
@@ II-B Assessing enjoyment
Enjoyment is generally evaluated through self-reported questionnaires tailored to the specific application domain. For instance, the Quality of Life Enjoyment and Satisfaction Questionnaire (Q-LES-Q) [31] and Physical Activity Enjoyment Scale (PACES) [32] are used in healthcare applications. In HCI and HRI research, enjoyment is not often the primary focus for evaluating user perceptions, but is typically included as part of a more comprehensive model [15]. Enjoyment frequently appears as a self-reported measure, either as a construct within established models like the Unified Theory of Acceptance and Use of Technology (UTAUT) [22], or as single-item measures in custom questionnaires (e.g.,“Did you feel fun?” [33], “Was playing with the robot enjoyable/not enjoyable?” [34]). Enjoyment was found to be highly correlated with ‘satisfying’, ‘entertaining’, ‘exciting’,‘fun’, and ‘interesting’ in HRI [35]. Technology Acceptance Model (TAM) [36] was also adapted to HRI by incorporating measures of affect and cognition to improve its accuracy in explaining technology adoption, and this adaptation included questions about perceived enjoyment [37].
@@
While user enjoyment is commonly measured through self-reporting in HRI, using it as a self-report metric has several limitations, such as conforming to perceived norms or researcher expectations or the (in)ability to recall the events and report correctly from memory [39, 16]. In addition, it is often desirable to estimate what a user is feeling by assessing it from an external perspective when self-reporting is not possible or the goal is to automate behavior at the dialogue exchange level during interactions in real-time. External assessments and self-reports are not mutually exclusive; rather, they can complement each other, with the results of one method potentially serving to validate the other. While prior research has relied on smiles and laughter to automatically classify user enjoyment (e.g., in storytelling contexts) [17], these signals can be ambiguous, contradictory, and highly context-dependent [40, 41]. Moreover, automatic classification of enjoyment would benefit from incorporating a broader range of multimodal cues, as enjoyment is often expressed through more than just facial expressions.
@@ II-C Assessing enjoyment in conversations
User satisfaction is often used in conversation research [42, 43, 13], but is typically evaluated in relation to a task as previously described in Section II-A [44, 45]. For instance, the Paradigm for Dialogue System Evaluation (PARADISE) [44] is a framework for evaluating user satisfaction in dialogue systems, based on self-reported satisfaction on a dialogue level and is influenced by other metrics, such as task success in travel booking and accessing emails. Similarly, interaction quality, which evaluates user satisfaction from an external perspective on the exchange level, was analyzed by three annotators in bus schedule inquiries with chatbots over phone calls from a data corpus of 200 dialogues and a lab study with 38 subjects [45]. An autonomous system was developed based on their ratings, which correlated highly with them, but not with users’ self-reported satisfaction scores, which was attributed to the subjectivity of the measure and variability in user perceptions. User satisfaction has also been measured in the text domain with an annotation protocol similar to our study, based on a dataset of 1000 dialogues between 50 users and a chatbot on attentive listening and conversations about animals [12]. Two annotators were recruited and an annotator instruction session was conducted. The annotators were requested to go through the conversation exchanges once, without going back or looking at the history. The annotators used three metrics of user satisfaction: ‘smoothness of the conversation’, ‘closeness perceived by the user towards the system’, and ‘willingness to continue the conversation’, rated from 1 to 7. However, no agreement was found between the annotators, even after changing the granularity of the scale to two levels, low and high, showing the complexity of evaluating a subjective measure from an external perspective. Another study by Wei et al [13] developed a multimodal model for user satisfaction based on conversations with a virtual agent. User satisfaction was annotated by the wizard controlling the agent, the users themselves, and external annotators who rated dialogue exchanges on metrics like topic continuance, external sentiment, and self-sentiment. Notably, two metrics directly linked satisfaction to enjoyment. The resulting multimodal model outperformed human annotators in evaluating user satisfaction at the overall interaction level.
@@
Our user enjoyment scale was developed based on the work of Reimnitz and Rauer [18], which is the only scale identified that specifically evaluates user enjoyment in conversations from an external perspective. The study assessed the enjoyment in conversations between 64 married couples and compared that to each spouse’s marital happiness. For measuring enjoyment, they developed a scale for observational coding that took into consideration affective signs and the tone of the interaction. The scale ranged from 1 (very low enjoyment) to 7 (very high enjoyment), with 3 as a neutral anchor. Two annotators took into account both affective signs (e.g., mutuality of the interaction, tone of voice, consistent mutual gaze, facial expressions, physical touching, body language) and the tone of the interaction (i.e., neutral, enthusiastic, and delightful) when rating enjoyment. The annotators had good to moderate agreement, with intraclass correlation (ICC) on 20% of the interactions. The study found that couples who displayed high enjoyment in their conversations also reported having a happier marriage. This aligns with prior research in human relationships, which found that mutually enjoyable behavior leads to increased intimacy, trust, security, and satisfaction in long-term relationships [46], signaling that enjoyment could be highly influential in achieving long-term HRI.
@@
Similar to this scale, our study analyzes open-domain, dyadic conversations outside of task-oriented settings, but with a focus on evaluating enjoyment in conversations with autonomous robots. Although prior research assessed autonomous conversational systems with related metrics that touch upon aspects of enjoyment, these metrics capture it only indirectly. Our study, instead, emphasizes understanding enjoyment within conversations with robots. To the best of our knowledge, no prior study measures enjoyment in such settings from an external perspective that takes into account multimodal aspects of HRI. Our work seeks to bridge this gap by proposing a scale that captures user enjoyment within conversation at both the exchange and overall interaction levels, with the potential to be used for autonomous systems to adapt the conversations on the fly.
III Data
As outlined in Section I, achieving and maintaining user enjoyment is important for encouraging continued interactions with robots, especially in daily encounters. This becomes particularly prominent for companion robots for older adults that aim to provide social and emotional support to mitigate loneliness in their daily lives. This work builds upon the data from [10] on the participatory design development (with two studies) of an autonomous companion robot that integrates an LLM for conversations with older adults, to build the user enjoyment scale for conversational HRI and evaluate enjoyment.
This section summarizes the robot architecture used in both studies, outlines the studies, and describes how the data was selected from those studies to create the scale (alignment) and evaluate it.